OpenStudy (anonymous):

find the slope of the line that goes through the given points (-5,8),(-2,-7

terenzreignz (terenzreignz):

Slope is all about differences: differences and their quotient. So... what are the differences between the y coordinates and the x coordinates?

terenzreignz (terenzreignz):

In maths terms, you get the slope \(\large \color{blue}m\) of the line through two points \(\large (x_1, y_1)\) ad \(\large (x_2,y_2)\) is given by \[\LARGE \color{blue}m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

i dont know the formula but i would minus the x's and y's together, -2--5=3 and -7-8=15 therefore 3 on the x axis and 15 on the y axis a gradient of 5?

OpenStudy (anonymous):

m=-7-8/-2-(-5) =-15/-2+5 =-15/3 m=-5